# Decentralized real-time iterations for distributed NMPC

### v2

@@ IV-A Outer Convergence
    q-linearly, if $\|p^{k+1}-p^{\star}\|\leq c\|p^{k}-p^{\star}\|$ for all $k\geq k_{0}$, for some $10<c<$ 1, and for some $k_{0}\in\mathbb{N}_{0}$.
@@
For all $i\in\mathcal{S}$, denote the set of active inequality constraints at $z_{i}^{\star}$ as $\mathcal{A}_{i}\doteq\{j\in\{1,\dots,n_{h_{i}}\}\;|\;[h_{i}(z_{i}^{\star})]_{j}=0\}$. Likewise, we denote the set of inactive inequality constraints as $\mathcal{I}_{i}\doteq\{j\in\{1,\dots,n_{h_{i}}\}\;|\;[h_{i}(z_{i}^{\star})]_{j}<0\}$. Recall that we assume the functions in NLP (5) to be twice continuously differentiable. An exact SQP scheme is locally convergent under the following assumption [Boggs1995].
@@
Lemma 2 (Exact SQP convergence [Boggs1995, Thm. 3.1])
@@
Assumption 6 (ii) is slightly stronger than the standard Second-Order Sufficient Condition (SOSC) [Boggs1995, A4], because we exclude the conditions $[{\nabla h_{i}(z_{i}^{\star})^{\top}]_{\mathcal{A}_{i}}z_{i}=0}$ and $E_{i}z_{i}=0$ on $z_{i}$ when demanding positive definiteness of the Hessian. Thus, QP (6) has a unique KKT point if $p^{k}\approx p^{\star}$, which is not necessarily the case under the standard SOSC condition [Robinson1974].
@@
Lemma 3 (Inexact SQP convergence)
@@
Hence, if $p^{0}\in\mathcal{B}(p^{\star},\varepsilon_{2})$ and if $\varepsilon_{2}\leq\text{min}(\varepsilon_{1},(\bar{a}_{p}-a)/((1+a)C))$, then the sequence $\{p^{k}\}$ converges q-linearly to $p^{\star}$. ∎
Lemma 3 guarantees local SQP convergence, if the inexact QP solutions satisfy inequality (9) as in [Zanelli2021b, Lem. 3.1.10]. As we show next, the solutions produced by ADMM meet this requirement.
IV-B Inner Convergence
@@
Lemma 4 (ADMM convergence)
Suppose Assumption 5 holds and let $p^{\star}$ denote a KKT point which satisfies Assumption 6. Form QP (6) at a primal-dual point $p^{k}=(z^{k},\nu^{k},\mu^{k},\lambda^{k})$ using the exact Hessian $H^{k}=\nabla_{zz}^{2}L(z^{k},\nu^{k},\mu^{k})$. Initialize Algorithm 1 with $z_{i}^{k}$ and $E_{i}^{\top}\lambda^{k}$ for all $i\in\mathcal{S}$. Then, there exist an ADMM contraction factor $10<a_{w}<$ 1 and constants $d_{2},\varepsilon>0$ such that the following holds for all $10<a<$ 1. If $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$ and if
@@
We can regard QP (6) formed at $p^{k}\approx p^{\star}$ as a perturbed version of the QP formed at $p^{\star}$. Because of Assumption 5, we can apply the BST [Fiacco1983, Thm. 3.2.2] to QP (6) and thus there exists $\varepsilon>0$ such that the solution $p^{k,\star}_{\mathrm{QP}}$ to the perturbed QP (6) is regular for all $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$. The stronger SOSC condition (12) also holds if $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$, which follows by adjusting the proof of the BST to the fact that Assumption 6 (ii) is slightly stronger than SOSC [Fiacco1983, Lem. 3.2.1].
(b) From (a) we have that LICQ holds at $p^{k,\star}_{\mathrm{QP}}$ for all $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$. Thus, the Jacobian $\nabla g_{i}^{k\top}\in\mathbb{R}^{n_{g_{i}}\times n_{i}}$ has full row rank for all $i\in\mathcal{S}$. Recall that for any $A\in\mathbb{R}^{m\times n}$, by the fundamental theorem of linear algebra, the null space of $A$ and the range space of $A^{\top}$ together form $\mathbb{R}^{n}$ [Nocedal2006, p. 603]. For all $i\in\mathcal{S}$, we can thus decompose any point $y_{i}\in\mathbb{R}^{n_{i}}$ via the null space method [Nocedal2006, Ch. 16] as
@@
$w^{l}-w^{k,\star}_{\mathrm{QP}}=\begin{bmatrix}I&0&0&0\\ 0&0&0&E^{\top}/\rho\end{bmatrix}\left(p^{l}-p^{k,\star}_{\mathrm{QP}}\right)$
@@
The algorithm is only well-defined if $l_{\mathrm{max}}\geq 1$ and thus we obtain the inner iteration bound (11). ∎
@@
Lemma 5 (Lipschitz constant of the subsystem QP)
@@
Furthermore, $\mu_{\mathcal{I}}^{l+1}=0$. The KKT matrix $K^{k}$ is invertible for all $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$, because of LICQ and the stronger SOSC condition (12), cf. [Nocedal2006, Lem. 16.1]. Rearranging thus yields
@@
Since $\mu_{\mathcal{I}}^{l+1}=\mu_{\mathcal{I},\mathrm{QP}}^{k,\star}=0$, we obtain (14) with $d_{2}$ given in (20). ∎
@@
Lemma 6 (ADMM contraction factor)
@@
$w^{l+1}-w^{k,\star}_{\mathrm{QP}}=A^{k}\left(w^{l}-w^{k,\star}_{\mathrm{QP}}\right).$ (22)
@@
for all $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$. Finally, we obtain the worst-case ADMM contraction factor $a_{w}$ inside $\mathcal{B}(p^{\star},\varepsilon)$ via (23). ∎
Remark 3 (Multiple ADMM iterations per QP)
@@
IV-C Optimizer Convergence with Limited Inner Iterations
@@
Theorem 1 (dSQP convergence)
Suppose Assumption 5 holds and let $p^{\star}$ denote a KKT point which satisfies Assumption 6. Form QP (6) with the exact Hessian $H^{k}=\nabla_{zz}^{2}L(z^{k},\nu^{k},\mu^{k})$. Let $10<a<\bar{a}_{p}<$ 1. Then there exists a constant $\varepsilon>0$ such that the following holds with $d_{2}$ and $a_{w}$ computed via (20) and (23).
If $p^{0}\in\mathcal{B}(p^{\star},\varepsilon)$ and if the number $l_{\mathrm{max}}$ of ADMM iterations per SQP iterations satisfies (11), then the sequence $\{p^{k}\}$ generated by Algorithm 2 converges to $p^{\star}$ and, for all $k\in\mathbb{N}_{0}$, satisfies $\|p^{k+1}-p^{\star}\|\leq\bar{a}_{p}\|p^{k}-p^{\star}\|.$square
Proof:
By the outer convergence result (Lemma 3), there exists $\varepsilon_{2}>0$ such that the following holds. If the ADMM solutions to QP (6) satisfy the sufficient accuracy requirement (9) for the inexact SQP scheme and if $p^{0}\in\mathcal{B}(p^{\star},\varepsilon_{2})$, then the sequence $\{p^{k}\}$ generated by Algorithm 2 converges q-linearly to $p^{\star}$. Let $\varepsilon\leq\varepsilon_{2}$. From the inner convergence (Lemma 4), there exists a radius $\varepsilon>0$ such that the ADMM solutions satisfy (9) for all $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$. Thus, we obtain q-linear convergence in the outer iterations. ∎
Remark 4 (Relation to [Stomberg2022a])
An earlier version of dSQP is presented in [Stomberg2022a]. There, an inexact Newton-type stopping criterion terminates ADMM dynamically in each SQP step and control applications were not considered. Theorem 1 guarantees convergence when the number of ADMM iterations per SQP iteration is fixed to $l_{\text{max}}$. This avoids the online communication of convergence flags and allows us to use dSQP in the proposed decentralized RTI scheme. Consequently, the convergence proof derived in this article differs from [Stomberg2022a] and, inspired by a centralized inexact SQP scheme from [Zanelli2021b], is centered around inequality (9) instead of inexact Newton methods. $\square$
V Decentralized Real-Time Iterations
@@
Consider the setpoint stabilization for the network $\mathcal{S}$ of dynamical systems with dynamics (1b). In each control sampling interval $\delta$, the state $x(t)$ is assumed to be measured and a constant input signal is applied to the system. We define the distributed NMPC control law $\kappa_{\mathrm{d}}:\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{p}}\rightarrow\mathbb{R}^{n_{u}}$ as the map from the centralized state $x(t)$ and the dSQP initialization $p^{0}(t)$ to the first part $\bar{u}^{k_{\text{max}}}(0)$ of the centralized input trajectory which is returned by Algorithm 2 with settings $k_{\text{max}}\in\mathbb{N}$ and $H_{i}=\nabla_{z_{i}z_{i}}^{2}L_{i}$. That is, in each NMPC step we apply dSQP with one or more outer iterations and with $l_{\text{max}}$ inner iterations per outer iteration to OCP (1), or respectively its NLP reformulation (5). Furthermore, we initialize dSQP with the OCP solution obtained in the previous NMPC step. Thus, the centralized system (2b) and dSQP form the closed-loop system optimizer dynamics (4), where $\Phi$ maps the current state $x(t)$ and dSQP output $p^{k_{\mathrm{max}}}(t)$ to the dSQP solution at the next time step. We are now ready to state our main result.
Theorem 2 (Decentralized RTI stability)
@@
Suppose that Assumptions 1 and 4 hold and that $\bar{p}(0)=0$. Further assume, for all $x\in\mathbb{X}_{\bar{V}}$, that Assumption 5 holds and that $\bar{p}(x)$ satisfies Assumption 6. Consider $a\in(0,1)$. For all $x\in\mathbb{X}_{\bar{V}}$, denote the dSQP convergence radius and contraction factor from Theorem 1 by $\varepsilon(x)$ and $\bar{a}_{p}(x)$, compute the constants $d_{2}(x)$ and $a_{w}(x)$ according to Lemmas 5 and 6, and assume the number of inner iterations $l_{\mathrm{max}}$ satisfies (11). Let $\hat{r}_{p}=\min_{x\in\mathbb{X}_{\bar{V}}}\varepsilon(x)$ and $a_{p}=\max_{x\in\mathbb{X}_{\bar{V}}}\bar{a}_{p}(x)$, compute $\bar{\delta}$ and $\tilde{r}_{p}$ via (29), and suppose the control sampling interval satisfies $\delta\leq\bar{\delta}$. Then, the following holds.
If the initial state $x(0)\in\mathbb{X}_{\bar{V}}$ and if Algorithm 2 is initialized with $p^{0}(0)\in\mathcal{B}(\bar{p}(x(0)),\tilde{r}_{p})$ in the first NMPC step, then the origin is a locally exponentially stable equilibrium of the closed-loop system-optimizer dynamics (4). $\square$
Proof:
@@
Consider the reformulation of OCP (2) as a partially separable NLP (5). According to the dSQP convergence Theorem 1, Algorithm 2 converges q-linearly to the primal-dual solution $\bar{p}(x(t))$ for any initialization $p^{0}(t)\in\mathcal{B}(\bar{p}(x(t)),\hat{r}_{p})$ and for all $x\in\mathbb{X}_{\bar{V}}$. Hence, dSQP satisfies the q-linear convergence Assumption 3. The Lipschitz continuity of the map $\bar{p}$ follows from Assumption 6 via the BST [Fiacco1983, Thm. 3.2.2], because the functions in NLP (5) are twice continuously differentiable and because $\nabla g$ is continously differentiable with respect to $x(t)$. That is, Assumption 2 holds. The exponential stability of the origin for the closed-loop system-optimizer dynamics therefore follows from Lemma 1. ∎
Remark 5 (Region of attraction)
The estimate region of attraction of the system-optimizer dynamics (4) is [Zanelli2021]
$\Sigma\doteq\left\{(x,p)\in\mathbb{R}^{n_{x}+n_{p}}\;|\;x\in\mathbb{X}_{\bar{V}},\;\|p-\bar{p}(x)\|\leq\tilde{r}_{p}\right\}.$
The set $\mathbb{X}_{\bar{V}}=\{x\in\mathbb{R}^{n_{x}}\;|\;V(x)\leq\bar{V}\}$ is the level set over which Assumptions 145, and 6 hold and $\tilde{r}_{p}$ can be computed via (29). $\square$
Remark 6 (Continuity of $V$, $\sqrt{V}$, and $\bar{p}$)
@@
Assumption 1 requires that $V$ is continuous for all $x\in\mathbb{X}_{\bar{V}}$ and that $\sqrt{V}$ is Lipschitz continuous at the origin. These conditions hold if $V$ is Lipschitz continuous over $\mathbb{X}_{\bar{V}}$ and twice continuously differentiable at the origin [Zanelli2021]. This is the case if Assumption 6 holds for all $x\in\mathbb{X}_{\bar{V}}$ [Fiacco1983, Thm. 3.4.1]. The Lipschitz continuity of $\bar{p}$ in Assumption 2 follows from the differentiability Assumption 5 and the regularity Assumption 6 via [Fiacco1983, Thm. 3.2.2]. $\square$
Remark 7 (Stability close to global optima)
@@
The careful reader will have noticed that the stability result of Theorem 2 requires the optimizer initialization to be close to the global minimum of the OCP. This assumption carries over from the centralized result [Zanelli2021, Thm. 25], which we invoke to prove stability. In practice, one will only obtain local minima for non-convex NLPs. However, we observe stability in simulations despite the convergence to local minima. To the best of our knowledge, centralized and decentralized RTI stability guarantees which only require optimizer intializations close to local minima are yet unavailable. $\square$
Remark 8 (Relation to [Hours2016])
@@
Real-time distributed NMPC is also addressed in [Hours2016] and we comment on similarities and differences to our approach. Both schemes share three important properties: First, they do not require a coordinator. Second, they apply an a-priori fixed number of optimizer iterations in each NMPC step to find a suboptimal control input. Third, the optimizer contraction for subsequent NMPC steps is proven via two key ingredients: (a) q-linear optimizer convergence for sufficiently many inner iterations, i.e., $l_{\text{max}}$ ADMM iterations per SQP step for dSQP and $M$ primal iterations for Algorithm 1 in [Hours2016], where $M$ is in the notation of [Hours2016], and (b), sufficient proximity between subsequent state measurements, i.e., by choosing the sampling interval sufficiently small or by direct assumption [Hours2016].
@@
On the other hand, the schemes differ in the employed optimization algorithm, in the order in which computations are executed on the subsystems, and in the obtained convergence result. The computationally expensive Step 3 in ADMM can be executed by all subsystems in parallel. In contrast, the decomposition scheme in [Hours2016] assigns subsystems into groups and the groups execute computation steps in sequence. And whereas Theorem 2 proves the local exponential stability of system and optimizer, [Hours2016, Thm. 5] bounds the optimizer suboptimality in closed-loop without addressing the system asymptotics. However, we invoke the more recent stability results from [Zanelli2021] to prove stability. Due to the q-linear optimizer convergence of Algorithm 1 of [Hours2016], we conjecture that Theorem 2 also holds under suitable assumptions if dSQP is replaced with the approach from [Hours2016]. $\square$
Remark 9 (Application to linear-quadratic DMPC)
@@
If NLP (5) is a convex QP, dSQP reduces to ADMM. In this case, the q-linear convergence of ADMM yields closed-loop stability for $l_{\mathrm{max}}\geq 1$ and sampling intervals below $\bar{\delta}$. This follows from Lemma 1 by replacing the convergence in $p$ in Assumption 3 with the ADMM convergence (13), cf. [Zanelli2021b]$\square$
V-A Implementation Aspects

### v3

@@ 4.1 Outer Convergence
    q-linearly, if $\|p^{k+1}-p^{\star}\|\leq c\|p^{k}-p^{\star}\|$ for all $k\geq k_{0}$, for some $0<c<1$, and for some $k_{0}\in\mathbb{N}_{0}$.
@@
For all $i\in\mathcal{S}$, denote the set of active inequality constraints at $z_{i}^{\star}$ as $\mathcal{A}_{i}\doteq\{j\in\{1,\dots,n_{h_{i}}\}\;|\;[h_{i}(z_{i}^{\star})]_{j}=0\}$. Likewise, we denote the set of inactive inequality constraints as $\mathcal{I}_{i}\doteq\{j\in\{1,\dots,n_{h_{i}}\}\;|\;[h_{i}(z_{i}^{\star})]_{j}<0\}$. Recall that we assume the functions in NLP (5) to be three times continuously differentiable. An exact SQP scheme is locally convergent under the following assumption [36].
@@
Lemma 4 (Exact SQP convergence [36, 37]).
@@
Under Assumptions 5 and 6, the SQP scheme considered in Lemma 4 identifies the correct active set if $p^{k}\approx p^{\star}$ [36, Sec. 5]. The above convergence result thus follows from [37, Thm. 18.4]. Note that Assumption 6 (ii) is slightly stronger than the standard Second-Order Sufficient Condition (SOSC) [36, A4], because we exclude the conditions $\nabla h_{i}(z_{i}^{\star})^{\top}]_{\mathcal{A}_{i}}z_{i}=0$} and $E_{i}z_{i}=0$ on $z_{i}$ when demanding positive definiteness of the Hessian. This facilitates the ADMM convergence analysis presented below and ensures that QP (6) has a unique KKT point if $p^{k}\approx p^{\star}$. Note that the latter does not necessarily hold under the standard SOSC condition, which can be addressed in the convergence analysis of SQP schemes by selecting $p^{k+1}$ as the KKT point of QP (6) which is closest to $p^{k}$, cf. [38].
@@
Lemma 5 (Inexact SQP convergence).
@@
Hence, if $p^{0}\in\mathcal{B}(p^{\star},\varepsilon_{2})$ and if $\varepsilon_{2}\leq\text{min}(\varepsilon_{1},(\bar{a}_{p}-a)/((1+a)C))$, then the sequence $\{p^{k}\}$ converges q-linearly to $p^{\star}$.
Lemma 5 guarantees local SQP convergence, if the inexact QP solutions satisfy inequality (9) as in [39, Lem. 3.1.10]. As we show next, the solutions produced by ADMM meet this requirement.
4.2 Inner Convergence
@@
Lemma 6 (ADMM convergence).
Suppose Assumption 5 holds and let $p^{\star}$ denote a KKT point which satisfies Assumption 6. Form QP (6) at a primal-dual point $p^{k}=(z^{k},\nu^{k},\mu^{k},\lambda^{k})$ using the exact Hessian $H^{k}=\nabla_{zz}^{2}L(z^{k},\nu^{k},\mu^{k})$. Initialize Algorithm 1 with $z_{i}^{k}$ and $E_{i}^{\top}\lambda^{k}$ for all $i\in\mathcal{S}$. Then, there exist an ADMM contraction factor $0<a_{w}<1$ and constants $d_{2},\varepsilon>0$ such that the following holds for all $0<a<1$. If $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$ and if
@@
We can regard QP (6) formed at $p^{k}\approx p^{\star}$ as a perturbed version of the QP formed at $p^{\star}$. Because of Assumption 5, we can apply the BST [40, Thm. 3.2.2] to QP (6) and thus there exists $\varepsilon>0$ such that the solution $p^{k,\star}_{\mathrm{QP}}$ to the perturbed QP (6) is regular for all $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$. The stronger SOSC condition (12) also holds if $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$, which follows by adjusting the proof of the BST to the fact that Assumption 6 (ii) is slightly stronger than SOSC [40, Lem. 3.2.1].
(b) From (a) we have that LICQ holds at $p^{k,\star}_{\mathrm{QP}}$ for all $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$. Thus, the Jacobian $\nabla g_{i}^{k\top}\in\mathbb{R}^{n_{g_{i}}\times n_{i}}$ has full row rank for all $i\in\mathcal{S}$. Recall that for any $A\in\mathbb{R}^{m\times n}$, by the fundamental theorem of linear algebra, the null space of $A$ and the range space of $A^{\top}$ together form $\mathbb{R}^{n}$ [37, p. 603]. For all $i\in\mathcal{S}$, we can thus decompose any point $y_{i}\in\mathbb{R}^{n_{i}}$ via the null space method [37, Ch. 16] as
@@
[…]
@@
The algorithm is only well-defined if $l_{\mathrm{max}}\geq 1$ and thus we obtain the inner iteration bound (11).
@@
Lemma 7 (Lipschitz constant of the subsystem QP).
@@
Furthermore, $\mu_{\mathcal{I}}^{l+1}=0$. The KKT matrix $K^{k}$ is invertible for all $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$, because of LICQ and the stronger SOSC condition (12), cf. [37, Lem. 16.1]. Rearranging thus yields
@@
Since $\mu_{\mathcal{I}}^{l+1}=\mu_{\mathcal{I},\mathrm{QP}}^{k,\star}=0$, we obtain (14) with $d_{2}$ given in (20).
@@
Lemma 8 (ADMM contraction factor).
@@
[…]
@@
for all $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$. Finally, we obtain the worst-case ADMM contraction factor $a_{w}$ inside $\mathcal{B}(p^{\star},\varepsilon)$ via (23).
Remark 9 (Multiple ADMM iterations per QP).
@@
4.3 Optimizer Convergence with Limited Inner Iterations
@@
Theorem 10 (dSQP convergence).
Suppose Assumption 5 holds and let $p^{\star}$ denote a KKT point which satisfies Assumption 6. Form QP (6) with the exact Hessian $H^{k}=\nabla_{zz}^{2}L(z^{k},\nu^{k},\mu^{k})$. Let $0<a<\bar{a}_{p}<1$. Then there exists a constant $\varepsilon>0$ such that the following holds with $d_{2}$ and $a_{w}$ computed via (20) and (23).
If $p^{0}\in\mathcal{B}(p^{\star},\varepsilon)$ and if the number $l_{\mathrm{max}}$ of ADMM iterations per SQP iterations satisfies (11), then the sequence $\{p^{k}\}$ generated by Algorithm 2 converges to $p^{\star}$ and, for all $k\in\mathbb{N}_{0}$, satisfies $\|p^{k+1}-p^{\star}\|\leq\bar{a}_{p}\|p^{k}-p^{\star}\|.$ $\square$
Proof 4.5.
By the outer convergence result (Lemma 5), there exists $\varepsilon_{2}>0$ such that the following holds. If the ADMM solutions to QP (6) satisfy the sufficient accuracy requirement (9) for the inexact SQP scheme and if $p^{0}\in\mathcal{B}(p^{\star},\varepsilon_{2})$, then the sequence $\{p^{k}\}$ generated by Algorithm 2 converges q-linearly to $p^{\star}$. Let $\varepsilon\leq\varepsilon_{2}$. From the inner convergence (Lemma 6), there exists a radius $\varepsilon>0$ such that the ADMM solutions satisfy (9) for all $p^{k}\in\mathcal{B}(p^{\star},\varepsilon)$. Thus, we obtain q-linear convergence in the outer iterations.
Remark 11 (Relation to [28]).
An earlier version of dSQP is presented in [28]. There, an inexact Newton-type stopping criterion terminates ADMM dynamically in each SQP step and control applications were not considered. Theorem 10 guarantees convergence when the number of ADMM iterations per SQP iteration is fixed to $l_{\text{max}}$. This avoids the online communication of convergence flags and allows us to use dSQP in the proposed decentralized RTI scheme. Consequently, the convergence proof derived in this article differs from [28] and, inspired by a centralized inexact SQP scheme from [39], is centered around inequality (9) instead of inexact Newton methods. $\square$
5 Decentralized Real-Time Iterations
@@
Consider the setpoint stabilization for the network $\mathcal{S}$ of dynamical systems with dynamics (1b). In each control sampling interval $\delta$, the state $x(t)$ is assumed to be measured and a constant input signal is applied to the system. We define the distributed NMPC control law $\kappa_{\mathrm{d}}:\mathbb{X}_{0}\times\mathbb{R}^{n_{p}}\rightarrow\mathbb{R}^{n_{u}}$ as the map from the centralized state $x(t)$ and the dSQP initialization $p^{0}(t)$ to the first part $\bar{u}^{k_{\text{max}}}(0)=M_{u,p}p^{k_{\mathrm{max}}}(t)$ of the centralized input trajectory which is returned by Algorithm 2 with settings $k_{\text{max}}\in\mathbb{N}$ and $H_{i}=\nabla_{z_{i}z_{i}}^{2}L_{i}$. That is, in each NMPC step we apply dSQP with one or more outer iterations and with $l_{\text{max}}$ inner iterations per outer iteration to OCP (1), or respectively its NLP reformulation (5). Furthermore, we initialize dSQP with the OCP solution obtained in the previous NMPC step, i.e. $p^{0}(t)=p^{k_{\mathrm{max}}}(t-1)$. Thus, the centralized system (2b) and dSQP form the closed-loop system optimizer dynamics (4), where $\Phi$ maps the current state $x(t)$ and dSQP output $p^{k_{\mathrm{max}}}(t)$ to the dSQP solution at the next time step. We are now ready to state our main result.
Theorem 12 (Decentralized RTI stability).
@@
Suppose that Assumptions 1 and 4 hold and that $\bar{p}(0)=0$. Further assume, for all $x\in\mathbb{X}_{\bar{V}}$, that Assumption 5 holds and that $\bar{p}(x)$ satisfies Assumption 6. Consider a constant $a\in(0,1)$ for Theorem 10. For all $x\in\mathbb{X}_{\bar{V}}$, denote the dSQP convergence radius and contraction factor from Theorem 10 by $\varepsilon(x)$ and $\bar{a}_{p}(x)$, compute the constants $d_{2}(x)$ and $a_{w}(x)$ according to Lemmas 7 and 8, and assume the number of inner iterations $l_{\mathrm{max}}$ satisfies (11). Let $\hat{r}_{p}=\min_{x\in\mathbb{X}_{\bar{V}}}\varepsilon(x)$, $a_{p}=\max_{x\in\mathbb{X}_{\bar{V}}}\bar{a}_{p}(x)$, and compute $\bar{\delta}$ and $\tilde{r}_{p}$ via (31). Then, the following holds.
If the control sampling interval satisfies $\delta\leq\bar{\delta}$, then the origin is a locally exponentially stable equilibrium with region of attraction
[…]
of the closed-loop system-optimizer dynamics (4). $\square$
Proof 5.1.
@@
Consider the reformulation of OCP (1) as a partially separable NLP (5). According to the dSQP convergence Theorem 10, Algorithm 2 converges q-linearly to the primal-dual solution $\bar{p}(x(t))$ for any initialization $p^{0}(t)\in\mathcal{B}(\bar{p}(x(t)),\hat{r}_{p})$ and for all $x\in\mathbb{X}_{\bar{V}}$. Hence, dSQP satisfies the q-linear convergence Assumption 3. The Lipschitz continuity of the map $\bar{p}$ follows from Assumption 6 via the BST [40, Thm. 3.2.2], because the functions in NLP (5) are twice continuously differentiable and because $\nabla g$ is continously differentiable with respect to $x(t)$. That is, Assumption 2 holds. The exponential stability of the origin for the closed-loop system-optimizer dynamics therefore follows from Lemma 1.
Remark 13 (Region of attraction).
If the initial state $x(0)$ and the iterate $p^{k_{\mathrm{max}}}(0)$ returned by dSQP in the first NMPC step lie inside the region of attraction, i.e. if $(x(0),p^{k_{\mathrm{max}}}(0))\in\Sigma$}, then the system-optimizer dynamics converge to the origin. The set $\mathbb{X}_{\bar{V}}=\{x\in\mathbb{R}^{n_{x}}\;|\;V(x)\leq\bar{V}\}$ is the level set over which Assumptions 145, and 6 hold and $\tilde{r}_{p}$ in the definition of $\Sigma$ can be computed via (31). $\square$
Remark 14 (Continuity of $V$, $\sqrt{V}$, and $\bar{p}$).
@@
Assumption 1 requires that $V$ is continuous for all $x\in\mathbb{X}_{\bar{V}}$ and that $\sqrt{V}$ is Lipschitz continuous at the origin. These conditions hold if $V$ is Lipschitz continuous over $\mathbb{X}_{\bar{V}}$ and twice continuously differentiable at the origin [7]. This is the case if Assumption 6 holds for all $x\in\mathbb{X}_{\bar{V}}$ [40, Thm. 3.4.1]. The Lipschitz continuity of $\bar{p}$ in Assumption 2 follows from the differentiability Assumption 5 and the regularity Assumption 6 via [40, Thm. 3.2.2]. $\square$
Remark 15 (Stability close to global optima).
@@
The careful reader will have noticed that the stability result of Theorem 12 requires the optimizer initialization to be close to a global minimum of the OCP. This assumption carries over from the centralized result [7, Thm. 25], which we invoke to prove stability. In practice, one will only obtain local minima for non-convex NLPs. However, we observe stability in simulations despite the convergence to local minima. A key challenge when guaranteeing stability of MPC under local minima is that the OCP value function $V$ is unknown. Suboptimal MPC addresses this challenge for OCPs with terminal constraints by exploiting the feasibility of local minima [42, Sec. 2.7]. However, RTI schemes typically favor rapid computation at the cost of asymptotic feasibility. Thus, and to the best of our knowledge, centralized and decentralized RTI stability guarantees which only require optimizer initializations close to local minima are yet unavailable. $\square$
Remark 16 (Relation to [19]).
@@
Real-time distributed NMPC is also addressed in [19] and we comment on similarities and differences to our approach. Both schemes share three important properties: First, they do not require a coordinator. Second, they apply an a-priori fixed number of optimizer iterations in each NMPC step to find a suboptimal control input. Third, the optimizer contraction for subsequent NMPC steps is proven via two key ingredients: (a) q-linear optimizer convergence for sufficiently many inner iterations, i.e., $l_{\text{max}}$ ADMM iterations per SQP step for dSQP and $M$ primal iterations for Algorithm 1 in [19], where $M$ is in the notation of [19], and (b), sufficient proximity between subsequent state measurements, i.e., by choosing the sampling interval sufficiently small or by direct assumption [19].
@@
On the other hand, the schemes differ in the employed optimization algorithm, in the order in which computations are executed on the subsystems, and in the obtained convergence result. The computationally expensive Step 3 in ADMM can be executed by all subsystems in parallel. In contrast, the decomposition scheme in [19] assigns subsystems into groups and the groups execute computation steps in sequence. And whereas Theorem 12 proves the local exponential stability of system and optimizer, [19, Thm. 5] bounds the optimizer suboptimality in closed-loop without addressing the system asymptotics. However, we invoke the more recent stability results from [7] to prove stability. Due to the q-linear optimizer convergence of Algorithm 1 of [19], we conjecture that Theorem 12 also holds under suitable assumptions if dSQP is replaced with the approach from [19]. $\square$
Remark 17 (Application to linear-quadratic DMPC).
@@
If NLP (5) is a convex QP, dSQP reduces to ADMM. In this case, the q-linear convergence of ADMM yields closed-loop stability for $l_{\mathrm{max}}\geq 1$ and sampling intervals below $\bar{\delta}$. This follows from Lemma 1 by replacing the convergence in $p$ in Assumption 3 with the ADMM convergence (13), cf. [39]. $\square$
5.1 Implementation Aspects